OpenStudy (anonymous):

Can somebody help me. Let the function h be defined by h(x) = 2 cos(10x) + 12. The maximum value of h is attained at which of the following values of x? A. pi over 5 B. pi over 10 C. pi over 15 D. pi over 20 E. pi over 40

OpenStudy (anonymous):

i know how it works in trigonometry

OpenStudy (kayders1997):

I would just do trial and error plug in the values they give you and see which one is largest

OpenStudy (anonymous):

max of cosine is 1 at \(\theta =0,2\pi,4\pi,...\)

OpenStudy (anonymous):

set \[10x=2\pi\] solve for \(x\) in one step